# Skykatana: a scalable framework to construct sky masks for the Vera Rubin Observatory and large astronomical surveys

### v1

@@ 3.2 Model and Fitting
When halos are included, the radius–magnitude relation shows three distinct regimes. For bright stars ($G\lesssim 13.6$) the radii decrease rapidly with magnitude, reflecting the shrinking extent of the saturated halo. For fainter stars ($G\gtrsim 15$) the halo is negligible and the radii follow a shallower trend. Between these limits the relation transitions smoothly. For each band, we therefore fitted three consecutive power laws of the form
@@
one in each magnitude interval, enforcing continuity at $G_{\mathrm{Gaia}}=13.6$ and $G_{\mathrm{Gaia}}=15.0$. To further isolate the halo signal, these fits were performed using only DP1 sources brighter than conservative magnitude limits: 23.75, 25.00, 25.00, 25.00, 24.20, and 23.00 in u, g, r, i, z, y, respectively. When the halo is removed, the radius–magnitude relation becomes smooth across the full magnitude range and is well described by a single power law per band. Tables 1 and 2 summarize the fitted parameters for the halo and no-halo cases, respectively, and Fig. 3 shows the corresponding fitted relations. In the y band, residual halo-induced overdensities persist even after quality cuts, preventing a reliable fit.
@@ 3.3 Results and Validation
Using Gaia and Tycho-2 stars over the full HSC footprint, coupon2018 derived empirical radius–magnitude relations based on the depletion of source density around saturated stars. Their adopted functional form consists of two exponential branches, which correspond to exclusion radii of $\sim 50$$60^{\prime\prime}$ for $G_{\mathrm{Gaia}}\simeq 10$ and $\sim 5^{\prime\prime}$ for $G_{\mathrm{Gaia}}\simeq 17$. . Our ECDFS relations show broadly consistent trends: the exponential slopes at the faint end are comparable to those of HSC within uncertainties, while the bright-end radii are slightly smaller, reflecting the fact that the ECDFS data are derived from a more limited one-square-degree field. Overall, both studies demonstrate that empirical bright-star masking based on density profiles yield robust, survey-specific prescriptions that balance completeness and area loss.
Finally, we emphasize that the ECDFS covers only $\simeq 1$ deg2 and contains relatively few bright stars. Consequently, the bright-end radii presented here should be used with caution. As Rubin accumulates more data, we plan to repeat this analysis over wider and deeper regions, which will provide more reliable calibrations of the bright-star exclusion radii, especially for the brightest stars where DP1 statistics are currently limited.
@@ 4.3 Masking bright stars
The effect of bright stars in HSC can be decomposed into circular shapes due to the bright halo of stars (STARS_REGIONS), and into rectangular boxes due to saturated bleed trails along the camera CCDs (BOX_STARS_REGIONS, see coupon2018 for their definition).
@@ 4.5 Adding a stage for WISE
To create a mask for WISE artifacts (difraction spikes, star halos, ghosts, etc.) we pixelize a catalog of objects severely affected by them. These were queried from the Reject table of the AllWISE database111irsa.ipac.caltech.edu/Missions/wise.html, by selecting objects either spurious or contaminated by those artifacts in W1, W2 or W3 bands. It is entirely possible to create a mask for moon trails (i.e. scattered light due to off-field lunar illumination) with the same procedure, but omit the steps for the sake of brevity. Below, we pixelize the artifacts (WISE_FILE) and regenerate the combined mask. The resulting MOC is visualized in Figure 10.
@@ 5 Application II: Rubin bright-star masks over the entire survey footprint
As a second application of Skyatana, we demonstrate how to construct bright-star exclusion masks within the Rubin Science Platform environment by querying and pixelating Gaia stars on-demand. This example (albeit with a much simpler input stage) is covered in more detail in this notebook. In order to extract the various regions defined in the Rubin baseline footprint, please follow this notebook and the steps here to ingest it as a pipeline stage.
@@ 5.2 Building and inspecting the mask
The code proceeds as described in Sect. 2.3: it converts the MOC of the search region into BF chunks, adds a pixel around borders, queries the Gaia HATS endpoint for each chunk, assigns radii using the DP1 calibrated function, and pixelizes each exclusion disc into a bit-packed boolean map. The partial maps are then streamed and merged into a single stage of (customizable) name starmask. The result is a mask consisting of $\sim$1.5 billion pixels of order 15 (6.4” pixel size) produced by pixelating about 100 million stars. The mask has a total area of  4960 deg2 that with the above configuration is computed in around 2 hours in the RSP.
@@ 6 Discussion
Skykatana demonstrates that large-scale angular masking can be achieved efficiently through sparse pixelization, streaming I/O, and modular composition. Its design deliberately prioritizes memory efficiency and scalability, allowing users to build multi-billion-pixel masks with limited hardware resources. The framework bridges the gap between polygon-based methods such as mangle and modern HEALPix-based pipelines by providing an intermediate high-level abstraction that is intuitive and efficient. By using the dual “coverage/sparse” structure of HealSparse and combining it with MOC-based visualization, the code achieves a balance between precision and usability for survey-scale data analysis.
Another strength of the framework worth highlighting is its on-demand catalog querying mechanism. Rather than relying on precomputed, monolithic bright-star masks, Skykatana builds exclusion maps dynamically over user-defined regions, exploiting hierarchical chunking to keep the I/O and memory footprint constant. This approach is particularly well suited to distributed platforms such as the Rubin Science Platform with LSDB/HATS environments, where catalogs are stored as HEALPix-partitioned parquet files.
@@
Compared to the HSC-SSP bright-star mask approach, Skykatana generalizes the concept from static survey products to a reproducible procedural pipeline. The resulting maps are automatically extensible to any region, filter, completeness threshold, or user-defined magnitude-radius relation. Relative to MOCpy, which focuses on set operations over MOCs, Skykatana operates at much higher pixel orders and retains explicit boolean semantics for every sparse pixel.
@@
Several simplifying assumptions remain to be addressed. Circular exclusion zones approximate stellar halos but cannot fully capture the complex anisotropy of diffraction spikes, ghosts, or disperse light. In addition, boolean masking inherently discards all affected pixels equally, whereas partial weighting schemes could preserve valuable information in low-contamination regions. Extending Skykatana to support non-boolean (integer or floating point) maps, e.g., weight, depth, or completeness fields, would broaden its scope to probabilistic selection functions and variable-coverage analyses.
@@
The ECDFS field observed by Rubin is by definition a region devoid of very bright stars, whereas other fields that are also part of DP1 have not been observed to 10-year depth. As a consequence, the radius–magnitude relation derived in Section 3 is not well constrained for stars brighter than $\sim$8-8.5 mag. As the Rubin survey progresses, we will recalibrate the relation to characterize better its effec
@@ 7 Summary and conclusions
These examples show that arcsecond resolution masks of up to billions of pixels can be generated within few hours with modest computing resources while maintaining full reproducibility and traceability of parameters. Skykatana visualization methods provide immediate diagnostic power for assessing the spatial impact of masking decisions. Looking ahead, Skykatana can serve as a foundational layer for next-generation survey pipelines that demand spatial reproducibility at scale. By extending the current boolean representation to include weighted and continuous fields, the framework could integrate seamlessly with probabilistic catalog analyzes, joint likelihood mapping, and end-to-end selection function modeling for Rubin and beyond.
@@ Acknowledgements
This project was made possible through a collaboration with the LINCC Frameworks software engineering team as part of the Frameworks Incubator Program, which is supported by the LSST Discovery Alliance (LSST-DA).
@@
This research has also made extensive use of the following software: Astropy (astropy), Matplotlib (matplotlib), Pandas (pandas), Seaborn (seaborn), Healpy (healpy), SciPy (scipy), NumPy (numpy), HEALSparse (hsppypi), MOCpy (boch2019), HATS/LSDB schemas (caplar2025), The Jupyter Notebook (jupyter-notebook), and TOPCAT (topcat).
@@
This research use data from the NSF-DOE Vera C. Rubin Observatory (2025); Legacy Survey of Space and Time Data Preview 1 (10.71929/rubin/2570308) and materials from NSF-DOE Vera C. Rubin Observatory (2025); Rubin Observatory LSST Tutorials (10.11578/rubin/dc.20250909.20).

### v2

@@ 3.2 Model and Fitting
When halos are included, the radius–magnitude relation shows three distinct regimes. For bright stars ($G\lesssim 13.6$) the radii decrease rapidly with magnitude, reflecting the shrinking extent of the saturated halo. For fainter stars ($G\gtrsim 15$) the halo is negligible and the radii follow a shallower trend. Between these limits, the relation transitions smoothly. For each band, we therefore fitted three consecutive power laws of the form
@@
one in each magnitude interval, enforcing the continuity at $G_{\mathrm{Gaia}}=13.6$ and $G_{\mathrm{Gaia}}=15.0$. In this empirical relation, $A_{i}$ serves as the normalization constant defining the scale of the exclusion zone, while $B_{i}$ determines the rate at which the radius shrinks with increasing stellar magnitude, physically capturing the steepness of the PSF wings and the resulting halo contamination profile in each filter. To further isolate the halo signal, these fits were performed using only DP1 sources brighter than the conservative magnitude limits: 23.75, 25.00, 25.00, 25.00, 24.20, and 23.00 in u, g, r, i, z, y, respectively. When the halo is removed, the radius–magnitude relation becomes smooth across the full magnitude range and is well described by a single power law per band. Tables 2 and 3 summarize the fitted parameters for the halo and no-halo cases, respectively, and Fig. 3 shows the corresponding fitted relations. In the y band, residual halo-induced overdensities persist even after quality cuts, preventing a reliable fit.
@@ 3.3 Results and Validation
Using Gaia and Tycho-2 stars over the full HSC footprint, Coupon et al. (2018) derived empirical radius–magnitude relations based on the depletion of source density around saturated stars. Their adopted functional form consists of two exponential branches, which correspond to exclusion radii of $\sim 50$$60^{\prime\prime}$ for $G_{\mathrm{Gaia}}\simeq 10$ and $\sim 5^{\prime\prime}$ for $G_{\mathrm{Gaia}}\simeq 17$. . Our ECDFS relations show broadly consistent trends: the exponential slopes at the faint end are comparable to those of HSC within uncertainties, while the bright-end radii are slightly smaller, reflecting the fact that the ECDFS data are derived from a more limited one-square-degree field. Overall, both studies demonstrate that empirical bright-star masking based on density profiles yield robust, survey-specific prescriptions that balance completeness and area loss.
We emphasize that the bright-star radius–magnitude relations derived in this section are based on a single $\sim 1\mathrm{deg}^{2}$ DP1 field (ECDFS), which provides limited statistics at the bright end and does not allow for a robust calibration in the $y$ band. Consequently, these relations should be interpreted as field-specific characterizations rather than universal prescriptions. Variations in survey depth, observing conditions, stellar density, and the effective area of the available fields are expected to introduce field-to-field differences. In particular, future analyses based on substantially larger sky areas will sample a broader population of bright stars and are therefore likely to yield different and more robust estimates of the characteristic exclusion radii.
@@ 4.3 Masking bright stars
The effect of bright stars in HSC can be decomposed into circular shapes due to the bright halo of stars (STARS_REGIONS), and into rectangular boxes due to saturated bleed trails along the camera CCDs (BOX_STARS_REGIONS, see Coupon et al. 2018 for their definition).
@@ 4.5 Adding a stage for WISE
To create a mask for WISE artifacts (difraction spikes, star halos, ghosts, etc.) we pixelize a catalog of objects severely affected by them. These were queried from the Reject table of the AllWISE database333irsa.ipac.caltech.edu/Missions/wise.html, by selecting objects either spurious or contaminated by those artifacts in W1, W2 or W3 bands. It is entirely possible to create a mask for moon trails (i.e. scattered light due to off-field lunar illumination) with the same procedure, but omit the steps for the sake of brevity. Below, we pixelize the artifacts (WISE_FILE) and regenerate the combined mask. The resulting MOC is visualized in Figure 10 and Table 6 lists the key configuration parameters adopted.
@@
Parameter Value HEALPix sparse order (footprint) 13 HEALPix sparse order (output) 15 Coverage order 4 Bit-packed storage Enabled Logical combination Footprint $\cap$ Patch - Stars - WISE artifacts
Table 6: Key configuration parameters for the HSC–WISE demonstration (Application I).
@@ 5 Application II: Rubin bright-star masks over the entire survey footprint
As a second application of Skykatana, we demonstrate how to construct bright-star exclusion masks within the notebook aspect of the Rubin Science Platform environment by querying and pixelating Gaia stars on-demand. This example (albeit with a much simpler input stage) is covered in more detail in this notebook. In order to extract the various regions defined in the Rubin baseline footprint, please follow this notebook and the steps here to ingest it as a pipeline stage.
@@ 5.2 Building and inspecting the mask
The code proceeds as described in Sect. 2.3: it converts the MOC of the search region into BF chunks, adds a pixel around borders, queries the Gaia HATS endpoint for each chunk, assigns radii using the DP1 calibrated function, and pixelizes each exclusion disc into a bit-packed boolean map. The partial maps are then streamed and merged into a single stage of (customizable) name starmask. The result is a mask consisting of $\sim$2 billion pixels of order 15 (6.4” pixel size) produced by pixelating about 110 million stars. The mask due to stars has a total area of  6480 deg2 that with the above configuration is computed in around 68 minutes in the RSP. Key parameters used in this application are listed in Table 7.
@@
Parameter Value Search stage WFD low-dust footprint HEALPix sparse order 15 HEALPix coverage order 5 Gaia magnitude range $9<G<20$ Chunk target area 800 deg2 Coarse BFS order 5 Milky Way exclusion $|b|<15^{\circ}$ + elliptical bulge mask Radius function Empirical DP1 $R(G)$ relation (85% completeness) Bit-packed storage Enabled Dask workers / memory 3 workers, 6 GB each (max)
Table 7: Key configuration parameters for the Rubin bright-star mask demonstration (Application II).
The on-demand star-mask construction described here leverages LSDB/HATS catalogs hosted within the Rubin Science Platform. However, the workflow is not RSP-specific. Users can reproduce the analysis outside of RSP by providing any local or remote catalog accessible via Dask, and just pointing the query interface to the HATS catalog folder. For environments without LSDB, Skykatana also supports building star masks from user-supplied catalogs in a variety of formats via the build_circ_mask() method, enabling full reproducibility on local computing systems.
In Table 8 we summarize the computational requirements for two representative Skykatana applications. The reported figures correspond to end-to-end wall-clock execution times, and peak memory usage. We note that the catalog-query component of the pipeline is executed on the Rubin Science Platform, which is deployed on a shared cloud infrastructure. As a result, absolute query times can vary between runs depending on system load, caching effects, and HATS endpoint location.
Application Area (deg2) Order $t_{wall}$ $t_{pix}$ $m_{peak}$ HSC–WISE (local) 69 13–15 6 s Rubin WFD (RSP) 17645 15 65 m 53 m 2.3 GB
Table 8: Summary of computational requirements for representative Skykatana applications. The columns correspond to the search stage area, wall clock time, pixelization time (single thread), and peak memory usage. The RSP example was run with 3 Dask workers and 1 thread per worker. For the HSC–WISE application we quote only the end-to-end runtime (cell 3 of the notebook), as the total execution time is sufficiently short that a robust separation between I/O and pixelization costs is not meaningful.
@@ 6 Discussion
Skykatana demonstrates that large-scale angular masking can be achieved efficiently through sparse pixelization, streaming I/O, and modular composition. Its design deliberately prioritizes memory efficiency and scalability, enabling users to build multi-billion-pixel masks within the constrained resources typical of shared-notebook and cloud-based environments. The framework bridges the gap between polygon-based approaches such as mangle and modern HEALPix-based pipelines by providing a high-level abstraction that is both intuitive and computationally efficient. By leveraging the dual “coverage/sparse” structure of HealSparse and combining it with MOC-based visualization, Skykatana achieves a balance between geometric precision, scalability, and usability for survey-scale data analysis.
Modern wide-area surveys such as Rubin, Euclid, and Roman increasingly require angular masks at resolutions corresponding to billions of HEALPix pixels, particularly to mitigate stellar contamination, artifacts, and spatially varying completeness. In practice, these masks must be constructed, inspected, and iterated upon within shared, notebook-based, or cloud environments where memory and I/O resources are constrained. Skykatana is explicitly designed for this regime, enabling the construction, manipulation, and interactive visualization of such masks through sparse, bit-packed representations and streaming I/O. A central contribution of the framework is its ability to construct masks directly from distributed catalog queries rather than from pre-materialized full-sky products, by streaming only the relevant sky regions and magnitude ranges through a chunked workflow. Although demonstrated here using the Rubin Science Platform infrastructure, the same design naturally applies to other large-survey analysis ecosystems based on distributed data access.
To represent survey masks on the sphere, we adopt the HEALPix tessellation. Although HEALPix pixels are not identical in shape, all pixels at a given resolution have exactly the same area and form a hierarchical and iso-latitude grid (Górski et al. 2005). This equal-area property is particularly relevant for angular masks and clustering analyses, as it avoids artificial density fluctuations introduced by variable pixel areas. The hierarchical structure further enables efficient multi-resolution representations, chunked I/O, and boolean mask algebra, while the iso-latitude sampling facilitates fast indexing and spherical operations. Alternative tessellations such as the Hierarchical Triangular Mesh (HTM; Kunszt et al. 2001) provide more regular geometric elements and fast indexing, but do not guarantee equal-area pixels and are less commonly adopted in modern survey masking frameworks. For these reasons, HEALPix has become the de facto standard in cosmological surveys (e.g., Planck, Gaia, Rubin, Euclid), and is adopted here for statistical robustness and interoperability.
@@
Compared to the HSC-SSP bright-star mask approach, Skykatana generalizes the concept from static survey products to a reproducible procedural pipeline. The resulting maps are automatically extensible to arbitrary sky regions, filters, completeness thresholds, and user-defined magnitude–radius relations. Relative to mocpy, which focuses on set operations over multi-order coverage maps, Skykatana operates at substantially higher pixel orders and retains explicit boolean semantics at the sparse-pixel level, enabling fine-grained control over mask geometry.
@@
Several simplifying assumptions remain to be addressed. Circular exclusion zones provide a practical approximation to stellar halos but cannot fully capture the complex anisotropy of diffraction spikes, ghosts, or diffuse scattered light. In addition, boolean masking inherently treats all affected pixels equally, whereas partial weighting schemes could preserve useful information in regions of low contamination. Extending Skykatana to support non-boolean (integer or floating-point) maps, such as weight, depth, or completeness fields, would broaden its scope to probabilistic selection functions and variable-coverage analyzes.
@@
Finally, the ECDFS field observed by Rubin is, by construction, a region largely devoid of very bright stars, whereas other fields included in DP1 have not yet been observed to full survey depth. As a consequence, the radius–magnitude relation derived in Section 3 is not well constrained for stars brighter than approximately 8–8.5 mag. As the Rubin survey progresses and deeper data become available over a wider range of environments, this relation will be recalibrated to better characterize the impact of bright stars on survey completeness.
@@ 7 Summary and conclusions
These examples show that arcsecond resolution masks of up to billions of pixels can be generated in a few dozen minutes with modest computing resources while maintaining full reproducibility and traceability of parameters. Skykatana visualization methods provide immediate diagnostic power for assessing the spatial impact of masking decisions. Looking ahead, Skykatana can serve as a foundational layer for next-generation survey pipelines that demand spatial reproducibility at scale. By extending the current boolean representation to include weighted and continuous fields, the framework could integrate seamlessly with probabilistic catalog analyzes, joint likelihood mapping, and end-to-end selection function modeling for Rubin and beyond.
In this work, the empirical demonstrations are intentionally focused on validating the performance, scalability, and reproducibility of the Skykatana framework using real survey data. The bright-star masking examples are presented as concrete use cases illustrating how catalog-driven masks can be constructed and applied at scale. Broader calibration validity, systematic field-to-field comparisons, and more complex masking schemes, such as anisotropic or environment-dependent masks, are not claimed here as validated results; rather, they are natural extensions enabled by the framework as larger and more diverse data sets become available. We are currently extending Skykatana to support floating-point (weighted) stellar masks that encode a probabilistic completeness correction rather than strict binary exclusion. Such mechanism can potentially retain a large fraction of area/sources that would otherwise get excluded, and even improve noise properties. A dedicated follow-up paper will present the method and quantify its impact on the large-scale structure (e.g. clustering measurements).
@@ Acknowledgements
The authors thank the referees for useful and constructive comments that improved this manuscript. This project was made possible through a collaboration with the LINCC Frameworks software engineering team as part of the Frameworks Incubator Program, which is supported by the LSST Discovery Alliance (LSST-DA).
@@
This research has also made extensive use of the following software: Astropy (Astropy Collaboration et al., 2013), Matplotlib (Hunter, 2007), Pandas (McKinney, 2010), Seaborn (Waskom, 2021), Healpy (Zonca et al., 2019), SciPy (Virtanen et al., 2020), NumPy (Harris et al., 2020), HEALSparse (Rykoff and Sánchez, 2023), MOCpy (Boch, 2019), HATS/LSDB schemas (Caplar et al., 2025), The Jupyter Notebook (Kluyver et al., 2016), and TOPCAT (Taylor, 2005).
@@
This research use data from the NSF-DOE Vera C. Rubin Observatory (2025); Legacy Survey of Space and Time Data Preview 1 (NSF-DOE Vera C. Rubin Observatory, 2025a) and materials from NSF-DOE Vera C. Rubin Observatory (2025); Rubin Observatory LSST Tutorials (NSF-DOE Vera C. Rubin Observatory, 2021).